### v1

# End-to-End Framework for Predicting the Remaining Useful Life of Lithium-Ion Batteries

Abstract
Accurate prediction of the Remaining Useful Life (RUL) is essential for enabling timely maintenance of lithium-ion batteries, impacting the operational efficiency of electric applications that rely on them. This paper proposes a RUL prediction approach that leverages data from recent charge-discharge cycles to estimate the number of remaining usable cycles. The approach introduces both a novel signal processing pipeline and a deep learning prediction model. In the signal preprocessing pipeline, a derived capacity feature $\dot{Q}(I,Q)$ is computed based on current and capacity signals. Alongside original capacity, voltage and current, these features are denoised and enhanced using statistical metrics and a delta-based method to capture differences between the current and previous cycles. In the prediction model, the processed features are then fed into a hybrid deep learning architecture composed of 1D Convolutional Neural Networks (CNN), Attentional Long Short-Term Memory (A-LSTM), and Ordinary Differential Equation-based LSTM (ODE-LSTM) modules. This architecture is designed to capture both local signal characteristics and long-range temporal dependencies while modeling the continuous-time dynamics of battery degradation. The model is further evaluated using transfer learning across different learning strategies and target data partitioning scenarios. Results indicate that the model maintains robust performance, even when fine-tuned on limited target data. Experimental results on two publicly available large-scale datasets demonstrate that the proposed method outperforms a baseline deep learning approach and machine learning techniques, achieving an RMSE of 101.59, highlighting its strong potential for real-world RUL prediction applications.
Keywords—Battery Health Management, Lithium-ion Batteries, Remaining Useful Life, Signal Processing, Deep Learning
[…]
[1]organization=AIWARE Limited Company, addressline=17 Huynh Man Dat Street, Hoa Cuong Bac Ward, Hai Chau District, city=Da Nang, postcode=550000, country=Vietnam
[…]
[2]organization=Faculty of Information Technology, Ton Duc Thang University, city=Ho Chi Minh City, postcode=700000, country=Vietnam
[…]
[3]organization=Software Engineering Department, FPT University, city=Da Nang, postcode=550000, country=Vietnam
1 Introduction
The growing global emphasis on sustainable energy has accelerated the advancement of battery technologies in recent years. Consequently, the lithium-ion battery (LIB) market is projected to surpass 170 billion USD by 2030 [26]. LIBs are widely used in applications such as grid energy storage systems, electric vehicles (EVs), and drones, all of which increasingly depend on effective Battery Health Management (BHM) systems [37, 5, 12] to ensure reliable performance. Typical BHM systems include key functionalities such as state of charge (SOC) estimation [28, 35], state of health (SOH) estimation [38, 20, 7], and remaining useful life (RUL) prediction [24, 23]. Among these, RUL prediction—which estimates the number of charge-discharge cycles remaining before a battery reaches its end-of-life (EOL)—is particularly crucial due to its implications for proactive maintenance, and safety assurance. During operation, LIBs undergo multiple forms of degradation during cycling, primarily caused by the loss of lithium inventory (LLI) and loss of active material (LAM) [25]. These challenges underscore the need for accurate RUL prediction models, not only to support real-time diagnostics and timely maintenance, but also to inform battery design [11].
In modern RUL prediction for LIBs, methods are mainly classified into two categories: curve-based and cycle-feature-based approaches. Curve-based methods model the degradation of a health indicator (e.g., capacity) over cycles using predefined functions (e.g., exponential decay), such as the double exponential model (DEM) [21], variational mode decomposition (VMD) [9], or empirical mode decomposition (EMD) [18, 15]. By fitting the historical data, these methods extrapolate the cycle at which the indicator falls below a predefined EOL threshold (e.g., 80% capacity [34]), thereby yielding the predicted number of remaining charge-discharge cycles (RUL). However, these methods have limited flexibility, as they assume a fixed degradation trend throughout the entire battery cycle life, which may not accurately capture real-world variability. Moreover, they often require a long degradation history to make accurate predictions. For instance, the method in [41] relies on data from the first 450 charge-discharge cycles to perform accurate RUL prediction, making it unsuitable for early-stage prognostics. In contrast, cycle-feature-based methods extract meaningful features from a few recent cycles [45, 22], enabling more adaptive, earlier predictions under different operating conditions across charge-discharge cycles.
In recent cycle-feature-based approaches, [45] predicts RUL of LIBs using only features extracted from the current charge-discharge cycle. The features including discharge capacity, initial battery capacity, and working condition-derived variables. However, this approach neglects the short-term degradation trends observable in the preceding cycles, which are critical for accurately capturing the battery’s dynamic aging behavior. Although [45] reports promising results, these outcomes are partially influenced by an imbalanced data split—70 batteries for training and only 7 for testing—which may lead to overfitting and reduced generalizability. In another study, [49] employs features extracted from the incremental capacity (IC) curve of a partial segment of the current charge-discharge cycle to predict the SOH and RUL. However, the method is evaluated only on a small battery dataset comprising just four battery cells discharged under constant current. This results in highly similar data distributions between the training and testing sets, making the prediction task easier and less representative of real-world scenarios. In contrast, a more robust and realistic approach is presented in [22], where features are extracted from 10 selected cycles within the 30 most recent charge-discharge cycles to predict RUL. This strategy effectively captures recent degradation patterns, enhancing predictive performance. Furthermore, [22] incorporates transfer learning to leverage degradation knowledge from a different source domain, improving the prediction accuracy in the target domain. Their method is validated on one of the largest publicly available battery datasets, comprising 55 cells for training and 22 cells for testing, with diverse discharge profiles to simulate real-world conditions.
In this work, we adopt the standard approach introduced in [22], where a window of sequential charge-discharge cycles is used as input to predict the number of remaining operational cycles. Our proposed method features a novel signal preprocessing pipeline and a hybrid deep learning (DL) prediction model. The signal preprocessing incorporates a newly derived feature based on interpolation; the gathered features then undergo denoising, feature extraction, and delta computation. The DL architecture comprises two parallel branches: a 1D Convolutional Neural Network (CNN) block for spatial feature extraction, and a recurrent branch that employs an Attentional Long Short-Term Memory (A-LSTM) network [16] to capture temporal dependencies, along with a Long Short-Term Memory Ordinary Differential Equation (LSTM-ODE) block to enhance dynamic sequence modeling capabilities. Both the CNN and LSTM-ODE blocks are custom-designed. We also evaluate the proposed hybrid DL model using transfer learning under various learning strategies and source data splitting scenarios. The results demonstrate that our model remains robust even when trained on limited source data. The model is validated on two of the most widely used lithium-ion battery datasets: one with varying charging profiles [34] and the other with diverse discharging profiles [22].
The remainder of this paper is organized as follows. Section 2 introduces the proposed method, including signal preprocessing, model architecture, and evaluation metrics. Section 3 describes the datasets and presents data analysis. Section 4 covers the experimental setup, model evaluations, and additional validation with transfer learning. Finally, Section 5 concludes the paper and outlines future work.
2 Proposed Method
The proposed method predicts RUL of lithium-ion batteries, defined as the number of charge-discharge cycles remaining from the current cycle $i$ until the battery reaches its EOL. The approach comprises two main stages: (1) a signal preprocessing pipeline for extracting informative features, and (2) a prediction model based on a hybrid deep learning architecture for estimating RUL. The overall architecture is illustrated in Figure 1. The details of these two components are presented in the following subsections.
Figure 1: Overall architecture of our proposed end-to-end framework.
2.1 Signal Preprocessing
We propose a signal preprocessing pipeline that efficiently constructs a feature tensor $\mathcal{F}^{(i)}_{30}$ for each input sample, representing a sequence of 10 consecutive charge-discharge cycles. The objective of the pipeline is to reduce noise and enhance the salient characteristics of the signals. The proposed preprocessing pipeline consists of the following steps:
  1. 1.
    Capacity Interpolation and Denoising: For each charge-discharge cycle $i$, we first generate evenly spaced values between the minimum and maximum of the current signal $I_{i}(t)$. A one-dimensional interpolation method [27] is then applied, using the sorted current values as the $x$-coordinates and the corresponding discharge capacities as the $y$-coordinates. This yields the derived capacity feature $\dot{Q}_{i}(Q,I)$, a technique we refer to as Capacity Derivative Tracking. Consequently, we obtain the data sequences $I_{i}(t)$, $V_{i}(t)$, $Q_{i}(t)$, and $\dot{Q}_{i}(Q,I)$. Each of these sequences is subsequently denoised using a selected filtering method—specifically, the Savitzky–Golay filter [33]—resulting in the smoothed signals $\tilde{x}_{i}\in\{\tilde{I}_{i}(t),\tilde{V}_{i}(t),\tilde{Q}_{i}(t),\tilde{\dot{Q}}_{i}(Q,I)\}$.
  2. 2.
    Statistical Feature Extraction: From each denoised sequence $\tilde{x}_{i}$, we compute a feature vector $F$ consisting of six statistical descriptors:
    $F(\tilde{x}_{i})=\left[\mu(\tilde{x}_{i}),\sigma(\tilde{x}_{i}),\min(\tilde{x}_{i}),\max(\tilde{x}_{i}),\mathrm{Var}(\tilde{x}_{i}),\mathrm{Median}(\tilde{x}_{i})\right],$ (1)
    where $\mu$ is the mean, $\sigma$ is the standard deviation, $\min$ and $\max$ denote the minimum and maximum values, $\mathrm{Var}$ is the variance, and $\mathrm{Median}$ is the median of the smoothed signal $\tilde{x}_{i}$.
  3. 3.
    Delta Feature Computation: To capture temporal dynamics, we compute the change in derived features $F(\tilde{\dot{Q}}_{i})$ over a cycle interval $\delta$ (set to 9 charge-discharge cycles, based on the optimal configuration identified in Figure 4):
    $\Delta F(\tilde{\dot{Q}}_{i})=F(\tilde{\dot{Q}}_{i})-F(\tilde{\dot{Q}}_{i-\delta}),$ (2)
    where $\Delta F(\tilde{\dot{Q}}_{i})$ represents the change in derived features between cycle $i$ and the previous cycle $i-\delta$.
  4. 4.
    Feature Fusion: For each charge-discharge cycle $i$ where $i>\delta$, we extract the following feature vectors: $F(\tilde{I}_{i})$, $F(\tilde{V}_{i})$, $F(\tilde{Q}_{i})$, $\Delta F(\tilde{\dot{Q}}_{i})$, each of size $\mathbb{R}^{6}$. Together, they form a fused feature matrix of size $\mathbb{R}^{4\times 6}$:
    $\mathcal{F}^{(i)}=\left[F(\tilde{I}_{i}),F(\tilde{V}_{i}),F(\tilde{Q}_{i}),\Delta F(\tilde{\dot{Q}}_{i})\right]\in\mathbb{R}^{4\times 6}.$ (3)
    For RUL prediction, each input sample comprises data from 10 uniformly spaced cycles selected within a 30-cycle window (i.e., one cycle every three cycles), following the strategy in [22]. The resulting input sample is:
    $\mathcal{F}^{(i)}_{30}\in\mathbb{R}^{10\times 4\times 6}.$ (4)
2.2 Prediction Model
Each input sample $\mathcal{F}^{(i)}_{30}$, as described in the previous section, is fed into the proposed prediction model to estimate the RUL. Our prediction model comprises three primary components: CNN block, A-LSTM block, and LSTM-ODE block. Among these, A-LSTM block is adopted from the prior work [16], while the other components are tailored.
In CNN block
The input tensor $\mathcal{F}^{(i)}_{30}$ is first reshaped into a sequence of 10 time steps, each represented by a 24-dimensional feature vector. This sequence is then passed through a stack of three one-dimensional convolutional layers with a shared kernel size of 5 and progressively increasing channel dimensions—$H$, $2H$, and $4H$, where $H=64$—consistent with the optimal configuration identified in Figure 5. Each convolution operation is followed by batch normalization [4], a Leaky ReLU activation [47], and dropout [42] for regularization. A fixed dropout rate of 0.3 is applied across all dropout layers in the prediction model. The resulting output is a feature map of shape $\mathbb{R}^{10\times 256}$, effectively capturing temporal dependencies and patterns within the cycle window.
A-LSTM Block
The feature map from the CNN block is fed into a two A-LSTM cells [16] with a sequence length of 10. The input size to the A-LSTM is $256$, and the hidden size is $128$. This block captures long-range temporal dependencies with attention mechanisms across the sequence. The hidden state at the final timestep is extracted and regularized via dropout, resulting in a output vector of shape $\mathbb{R}^{128}$.
ODE-LSTM Block
This block processes the input sample $\mathcal{F}^{(i)}_{30}$ by reshaping it into a matrix of shape $\mathbb{R}^{10\times 24}$, while preserving the temporal order. A continuous time sequence $t\in[0,1]$ with 10 uniformly spaced points is generated to simulate continuous-time dynamics. Both the reshaped input and time sequence are fed into two stacked ODE-LSTM cells, with each input size 24 and hidden size 256. The final hidden state is extracted and passed through a dropout layer, resulting in an output vector of size $\mathbb{R}^{256}$.
Our ODE-LSTM network extends the traditional long short-term memory (LSTM) [36] architecture by integrating a neural ordinary differential equation (ODE) to model the continuous evolution of hidden states between discrete inputs. The computation proceeds in two stages: (1) continuous-time evolution of the hidden state using an ODE solver, and (2) discrete-time update using the LSTM cell, as follows:
$\tilde{h}_{t}$ $=h_{t-1}+\text{ODE}_{\theta}(h_{t-1},t-1,t),$
$i_{t}$ $=\sigma(W_{i}[x_{t};\tilde{h}_{t}]),$
$f_{t}$ $=\sigma(W_{f}[x_{t};\tilde{h}_{t}]),$
$o_{t}$ $=\sigma(W_{o}[x_{t};\tilde{h}_{t}]),$
$\tilde{c}_{t}$ $=\tanh(W_{c}[x_{t};\tilde{h}_{t}]),$
$c_{t}$ $=f_{t}\odot c_{t-1}+i_{t}\odot\tilde{c}_{t},$
$h_{t}$ $=o_{t}\odot\tanh(c_{t}),$
where $\text{ODE}_{\theta}$ models the continuous transition of hidden states and is implemented using Euler integration:
[…]
Here, $f_{\theta}$ is a neural network (implemented as a linear layer), and [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes vector concatenation. $x_{t}$ is the input vector at time step $t$. $h_{t}$ is the hidden state at time $t$. $\tilde{h}_{t}$ is the hidden state after ODE evolution from $t_{t-1}$ to $t_{t}$. $c_{t}$ is the cell state at time $t$. $\sigma(\cdot)$ denotes the sigmoid activation function, and $\tanh(\cdot)$ denotes the hyperbolic tangent activation. $W_{i}$, $W_{f}$, $W_{o}$, and $W_{c}$ are trainable weight matrices for the input, forget, output, and candidate gates, respectively. $\odot$ denotes element-wise multiplication.
The output features from the CNN + A-LSTM and ODE-LSTM branches are concatenated into a single vector of dimension 384 and passed through a fully connected layer followed by a sigmoid activation to predict the RUL. Our proposed hybrid deep learning model is trained using Mean Squared Error (MSE) loss function, optimized with the AdamW optimizer [19]. The MSE loss is defined as:
$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}\left(y_{i}-\hat{y}_{i}\right)^{2}.$
Where $N$ is the number of training samples, $y_{i}$ denotes the ground truth RUL of the $i$-th cycling sample, and $\hat{y}_{i}$ denotes the predicted RUL of the $i$-th sample.
2.3 Evaluation Metrics
To assess the performance of RUL prediction, we employ three evaluation metrics: Root Mean Square Error ($RMSE$) [43], R-squared ($R^{2}$) [41], and Mean Absolute Percentage Error ($MAPE$) [6]. These metrics are defined as follows:
$RMSE(y_{i},\hat{y}_{i})=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}},$
$MAPE(y_{i},\hat{y}_{i})=\frac{1}{n}\sum_{i=1}^{n}\frac{|y_{i}-\hat{y}_{i}|}{y}\times 100,$
$R^{2}(y,\hat{y})=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}.$
Here, $y_{i}$ represents the observed RUL, and $\hat{y}_{i}$ represents the predicted RUL. Additionally, $y$ refers to the cycle life. The performance improves with lower values of $RMSE$ and $MAPE$, and a higher value of $R^{2}$.
3 Dataset
3.1 Dataset Description
3.1.1 First dataset
The dataset [34] features a comprehensive study of 124 lithium-ion batteries (LIBs) with LFP/graphite chemistry, focusing on various fast-charging conditions while maintaining consistent discharging parameters. Each battery in the dataset has a nominal capacity of 1.1 Ah and a nominal voltage of 3.3 V, with cycle lifespans ranging from 150 to 2,300 cycles, highlighting a broad range of battery longevity. All LIBs underwent uniform discharge procedures, involving a constant current discharge at a rate of 4 C until the voltage reached 2 V, followed by a constant voltage discharge at 2 V until the current dropped to C/50. Charging was conducted at rates between 3.6 C and 6 C under controlled temperature conditions of 30°C in an environmental chamber. The dataset comprises roughly 96,700 cycles, making it one of the largest datasets exploring various fast-charging protocols. It is divided into three sections: a training set with 41 LIBs, a validation set with 43 LIBs, and a test set with 40 LIBs.
3.1.2 Second dataset
The dataset [22], which involved varying discharge conditions while maintaining consistent fast-charging conditions, was collected from a battery degradation experiment. This experiment utilized 77 cells (LFP/graphite A123 APR18650M1A), each with a nominal capacity of 1.1 Ah and a nominal voltage of 3.3 V. Each cell underwent a distinct multi-stage discharge protocol, while all cells shared the same fast-charging protocol. The experiment was carried out in two thermostatic chambers, with a controlled temperature of 30°C. The dataset contains a total of 146,122 discharge cycles, making it one of the largest datasets that incorporate various discharge protocols. The cycle life of the cells ranges from 1,100 to 2,700 cycles, with an average of 1,898 cycles and a standard deviation of 387 cycles. The discharge capacity as a function of cycle number shows a broad distribution of cycle lives. The dataset is divided into two parts: In 55 LIBs predefined for training, 90% is used for the training set, 10% for the validation set, and the remaining 22 LIBs are used as the test set.
3.2 Data Analysis
Figure 2: Smoothed incremental capacity (dQ/dV) curves during charging for selected cycles on a battery cell from the second dataset: full voltage range (left) and zoomed-in region (3.4–3.6 V, right).
The degradation of LIBs through charge-discharge cycles is primarily caused by the LLI, leading to irreversible capacity fade. IC analysis, which plots $dQ/dV$ versus voltage, is a widely used diagnostic tool for identifying LLI-related degradation [1].
As shown in Figure 2, smoothed IC curves during constant current charging illustrate degradation trends in lithium-ion batteries. The position, intensity, and shape of $dQ/dV$ peaks provide key insights into electrochemical aging. Notably, the graphite peak associated with the $\mathrm{LiC_{6}\rightarrow LiC_{12}}$ phase transition [17] shifts left by over 10mV between Cycle 1 and Cycle 1501 (right plot), indicating lithium loss due to solid electrolyte interphase (SEI) formation. This irreversible process consumes lithium ions. Concurrently, the peak height drops by about 35%, signaling a decline in active lithium available for intercalation [29, 46].
Overall, LLI accounts for approximately 70% of the total observed capacity fade, based on the analysis of smoothed IC curves in Figure 2 using the methodology from [29]. Additionally, the figure reveals a capacity reduction of around 20% after 1500 cycles, with the degradation rate varying across the charge-discharge lifespan. These findings highlight the need for a robust RUL prediction method capable of capturing non-uniform aging patterns to reliably predict the RUL of lithium-ion batteries.
4 Experiments and Discussion
4.1 Experimental Setup
Our proposed RUL model is developed using the PyTorch framework. All experiments are performed on an NVIDIA 4090 GPU with 24GB of memory. Each experiment is trained for 10 epochs with a batch size of 128 and a constant learning rate of 0.0005. To minimize variability in the training process, each experiment is repeated 10 times, and the final prediction is calculated as the average of these multiple runs.
4.2 Evaluation of Signal Preprocessing
In the denoising step, to mitigate noise in voltage, current and capacity signals, four denoising techniques including Gaussian smoothing [44], Savitzky-Golay filtering [13], wavelet decomposition [8], and Kalman filtering-along with a no-denoising baseline (raw data), were evaluated on the second dataset. As shown in Figure 3, the Savitzky-Golay filter [32] achieved the optimal performance, yielding an RMSE of 169.93 (cycles). Compared to Gaussian smoothing and wavelet decomposition, the Savitzky-Golay method demonstrates superior capability in peak preservation and adaptive smoothing.
One key factor influencing the effectiveness of the Savitzky-Golay filter is the window length, which balances noise reduction and signal distortion. To determine the optimal configuration, various window lengths were tested. As shown in the right panel of Figure 3, a window length of 191 points provides the best trade-off between denoising and feature preservation. Consequently, the Savitzky-Golay filter with a window length of 191 is adopted in the denoising stage of our signal preprocessing pipeline.
Figure 3: RMSE comparison of different denoising methods (left) and the impact of the window length parameter on Savitzky-Golay method (right).
Figure 4 on the left shows that using only derived features results in the highest RMSE of 252.52, indicating that relying solely on derived features leads to poor predictive performance. In contrast, the feature combination of I, V, Q, and dQ(I) achieves the lowest RMSE value of 162.37, outperforming other feature sets. The strong contribution of dQ(I) highlights the effectiveness of the proposed Capacity Derivative Tracking method in estimating discharge data. By aligning discharge capacity values with corresponding current measurements, this method provides a more consistent representation of the current–discharge capacity relationship over time, enhancing the prediction performance of the model.
After the interpolation step, the difference in extracted features between each cycle and the one occurring delta steps earlier is computed. The choice of the delta parameter is important, as it determines the temporal resolution of feature variation. To assess its impact, we evaluated delta values ranging from 1 to 10, as shown in Figure 4 on the right. Among these, a delta value of 9 achieved the best performance, yielding the lowest RMSE of 162.37.
Figure 4: RMSE comparison of feature combinations (left) with delta set to 9, and the effect of delta on predictive performance (right).
4.3 Validation of Prediction Model
As shown in Table 1, the proposed signal preprocessing pipeline, which includes the derived capacity feature $\dot{Q}_{i}(Q,I)$, consistently improves the performance of all evaluated models. Notably, our proposed prediction model, CNN + A-LSTM + ODE-LSTM, achieves the best results, with the lowest RMSE of 158.09, highest $R^{2}$ score of 0.89, and lowest MAPE of 6.95%. These results confirm the effectiveness of combining CNN with A-LSTM and ODE-LSTM components, especially when enriched with the derived capacity feature.
Table 1: Performance comparison of deep learning models on the testing data of the second dataset.
Our Signal Preprocessing Model RMSE $\downarrow$ $\bm{R}^{2}$uparrow MAPE (%) $\downarrow$
W/o $\dot{Q}_{i}(Q,I)$ Attentional LSTM (A-LSTM) 174.83 0.84 8.14
CNN + A-LSTM 167.37 0.85 7.62
CNN + A-LSTM + ODE-A-LSTM 168.51 0.84 8.01
CNN + A-LSTM + ODE-LSTM (Ours) 163,31 0,85 7,54
With $\dot{Q}_{i}(Q,I)$ Attentional LSTM (A-LSTM) 168.09 0.87 7.47
CNN + A-LSTM 169.02 0.87 7.40
CNN + A-LSTM + ODE-A-LSTM 170.69 0.87 8.03
CNN + A-LSTM + ODE-LSTM (Ours) 158.09 0.89 6.95
To assess the impact of various configurations on the prediction model performance, we conducted a series of experiments evaluating RMSE across different settings. These included activation functions (GeLU [14], HardSwish [2], ReLU [3], LeakyReLU), ODE integration schemes (RK4 [39], Midpoint [48], Heun2/3 [30], Euler [31]), convolutional kernel sizes, hidden layer sizes, and the selection of the relative time variable $R_{t}$ (RT), as defined in the A-LSTM paper [16]. Each configuration was isolated while keeping other components fixed to ensure fair comparison. The results, visualized in Figure 5, indicate that specific choices—such as using LeakyReLU as the activation function, Euler for ODE integration, a kernel size of $k=5$, a hidden size of 64, and first RT in A-LSTM block yielded lower RMSE values. These optimal settings were highlighted in the figure to emphasize their contribution to improved predictive accuracy. This analysis validates the importance of fine-tuning model configurations to achieve optimal performance in the target task.
Figure 5: RMSE Comparison Across Configurations of the Proposed Prediction Model.
Figure 6 illustrates the RUL prediction performance of the proposed model across eight test battery cells taken from the test sets of two battery datasets. The red curves represent the predicted RUL, while the black curves denote the ground truth. Across all battery cells, the predicted RUL closely follows the actual degradation trend, confirming the model’s ability to generalize across diverse data distributions. Notably, the model successfully captures the nonlinear degradation behavior, particularly in cells such as b1c31 and 10-4, where the actual RUL exhibits mild fluctuations. Although slight overestimations appear in the early cycles and near EOL periods, the predictions remain consistently aligned with observed values.
Figure 6: Comparison of predicted RUL and actual RUL for some batteries from the first and the second datasets.
Table 2 presents a performance comparison between the proposed model and several baseline methods, including traditional machine learning models and the deep learning-based approach proposed in [22], across two battery datasets. On the first dataset, the proposed model significantly outperforms both ElasticNet [10] and XGBoostRegression [40], achieving the lowest RMSE of 101.59, the highest $R^{2}$ score of 0.82, and a substantially reduced MAPE of 8.29%. Similarly, on the second dataset, the proposed model surpasses all other methods, including the baseline model from [22], with an RMSE of 158.09, an $R^{2}$ score of 0.89, and the lowest MAPE of 6.95%. These results highlight the effectiveness and generalizability of the proposed approach across different data distributions.
Table 2: Comparison of the proposed approach with machine learning and existing approaches on both the first and second datasets.
Dataset Method RMSE $\downarrow$ $\bm{R}^{2}$uparrow MAPE (%) $\downarrow$
First dataset ElasticNet 319.98 -0.20 26.70
XGBoostRegression 156.33 0.70 12.19
Our model 101.59 0.82 8.29
Second dataset ElasticNet 575.54 -0.20 26.67
XGBoostRegression 169.88 0.87 7.37
[22] 186 0.804 8.72
Our model 158.09 0.89 6.95
4.4 Extended Validation of Transfer Learning on the Prediction Model
To evaluate the generalization capability of the proposed model across different battery datasets, we conducted a series of transfer learning experiments as detailed in Table 3. In cases involving upper and lower half splits, the training data were divided into two groups of battery cells based on cycle life: cells with a cycle life below the median of the cycle lives of the training set were assigned to the lower half, while those above the median were placed in the upper half. These experiments explore various scenarios in which the model is pre-trained on one dataset and fine-tuned on another, simulating real-world applications where labeled data in the target domain is scarce.
Table 3: Configuration of transfer learning strategies for each case, showing the datasets used for pre-training and fine-tuning with their respective testing sets.
Testing data of Case Pre-Training Fine-tuning
First dataset 1 Entire training data of the first dataset
2 Entire training data of the second dataset
3.1 Entire training data of the second dataset Upper half of training data from the first dataset
3.2 Entire training data of the second dataset Lower half of training data from the first dataset
4 Entire training data of the second dataset Entire training data of the first dataset
Second dataset 5 Entire training data of the second dataset
6 Entire training data of the first dataset
7.1 Entire training data of the first dataset Upper half of training data from the second dataset
7.2 Entire training data of the first dataset Lower half of training data from the second dataset
8 Entire training data of the first dataset Entire training data of the second dataset
First, we use Case 8 to investigate the impact of different block-freezing strategies during transfer learning, shown in Table 4. Notably, freezing only the A-LSTM block yields the best overall performance, suggesting that this component captures more transferable temporal features. In contrast, freezing all three blocks (CNN, A-LSTM, and ODE-LSTM) results in severe performance degradation, underscoring the importance of model flexibility during fine-tuning. Based on these findings, we adopt this block-freezing strategy of freezing only A-LSTM block for the next experiment, as shown in Table 5.
Table 4: Performance comparison under various block-freezing strategies applied durinng transfer learning in Case 8 (see Table 3).
Transfer Learning Strategy RMSE $\downarrow$ $\bm{R}^{2}$uparrow MAPE (%) $\downarrow$
Freezing CNN block (1) 166.18 0.88 7.21
Freezing A-LSTM block (2) 157.92 0.88 6.88
Freezing ODE-LSTM block (3) 159.64 0.87 6.79
(1) + (2) 170.05 0.86 7.48
(1) + (3) 179.4 0.85 8.04
(2) + (3) 161.98 0.86 7.01
(1) + (2) + (3) 282.69 0.69 12.79
The results, summarized in Table 5, demonstrate that direct transfer (Cases 2 and 6) without fine-tuning leads to poor performance, with high RMSE and negative $R^{2}$, indicating a significant mismatch in data distributions between the two datasets. However, introducing fine-tuning with partial data from the target domain (Cases 3.1, 3.2, 7.1, and 7.2) noticeably improves model accuracy, with Case 7.2 achieving a strong balance (RMSE = 201.63, $R^{2}=0.81$, MAPE = 8.63%). Full fine-tuning on the target domain after pre-training (Cases 4 and 8) yields performance close to or even better than training directly on the target dataset (Cases 1 and 5), with Case 8 achieving the best results on the second dataset (RMSE = 157.92, $R^{2}=0.88$, MAPE = 6.88%).
Table 5: Performance comparison of our proposed model on different training strategies with Transfer Learning on both first and second dataset.
Case RMSE $\downarrow$ $\bm{R}^{2}$uparrow MAPE (%) $\downarrow$
1 101.59 0.82 8.29
2 543.56 -2.58 46.66
3.1 160.98 0.61 14.4
3.2 290.75 0.00 23.31
4 123.44 0.76 10.64
5 158.09 0.89 6.95
6 800.23 -1.21 35.76
7.1 231.12 0.73 11.69
7.2 201.63 0.81 8.63
8 157.92 0.88 6.88
5 Conclusion
This study presents an robust approach for predicting the RUL of lithium-ion batteries by integrating an advanced signal processing pipeline with a deep learning prediction model. The signal preprocessing pipeline, comprising Capacity Interpolation and Denoising, Statistical Feature Extraction, Delta Feature Computation, and Feature Fusion, effectively captures the degradation dynamics over charge-discharge cycles. The proposed prediction model, combining CNN, A-LSTM, and ODE-LSTM blocks, demonstrates strong capabilities in modeling both discrete and continuous time behaviors of battery aging. Notably, the integration of our proposed CNN and ODE-LSTM block enhances the prediction model’s ability. The model is further evaluated using transfer learning across different learning strategies and source data partitioning scenarios. Results indicate that the model maintains robust performance even when trained on limited data, underscoring its adaptability and effectiveness in data-constrained environments. Experimental validation on two large-scale datasets confirms that our method significantly outperforms a baseline deep learning approach and traditional machine learning techniques. These results highlight the potential of the proposed approach for practical deployment in predictive maintenance systems, contributing to improved reliability and longevity of battery-powered applications. Future work will explore the generalizability of the method across different battery chemistries and usage scenarios. Additionally, efforts will be directed toward real-time deployment on edge devices.
Contact Information
For access to the code and further information about this proposed system, please contact AIWARE Limited Company at: https://aiware.website/Contact